what is n+n/3=600? Show your work please!
Answer by checkley71(8403)

n+n/3=600 combine the n terms by finding a common denominator (3).
(3n+n)/3=600 cross multiply.
(3n+n)=3*600
4n=1800
n=1800/4
n=450 answer.
proof:
450+450/3=600
450+150=600
600=600